OpenStudy (anonymous):

lim cot(x) x->pi from the right (the limit of cotx as x approaches pi from the right

rvc (rvc):

what is cot(pi) ? @maggielucci

OpenStudy (anonymous):

Remember the equation: \[\cot x = \frac{ \cos x }{ \sin x }\] which is continuos @ \[x = \frac{ \pi }{2 }\] SO , \[\lim_{x-> \frac{ \pi }{ 2 }} \cot x = \lim_{x-> \frac{ \pi }{ 2 }} \frac{ \cos x }{ \sin x} = \frac{ 0 }{ 1 } = 0 \] 0 or infinity

rvc (rvc):

good @Adjax

rvc (rvc):

it's only pi @Adjax

OpenStudy (anonymous):

so it would be positive infinity